OpenStudy (anonymous):

I will fanned and give you a medal !! 1) Evaluate the summation of 20 times 0.5 to the n minus 1 power, from n equals 3 to 12..

OpenStudy (tkhunny):

This? \(\sum\limits_{n=3}^{12}\left(20\cdot 0.5^{n-1}\right)\)

OpenStudy (anonymous):

Give me a sec to draw it

OpenStudy (anonymous):

\[\sum_{n=3}^{12?} 20(0.5)^{n-1}\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Same thing. Well, add them up! \(20\cdot\left(1/4 + 1/8 + 1/16 + ... + 1/2048\right)\) What do you get?

OpenStudy (anonymous):

Oh sorry, and do i just add straight across?

OpenStudy (anonymous):

the summation is a series of sums. The n starts at value 3 and each repetition of the sum the n increases with 1. You repeat the sums until n=12. That is what tkhunny did for you. But I'm writing it down into a broken down scheme to help you visualize what you're supposed to do mechanically \[n=3 \rightarrow 20\times0.5^{3} = ?\] \[n=4 \rightarrow 20\times0.5^{4} = ?\] \[n=5 \rightarrow 20\times0.5^{5} = ?\] ... \[n=12 \rightarrow 20\times0.5^{12} = ?\] \[\sum_{n=3}^{12}\] then means you add every separate result. So the total sum is the evaluation of the series you posted.

OpenStudy (tkhunny):

5⋅(1 +1/2+1/4+...+1/512) -- A little easier.

OpenStudy (anonymous):

Want me to solve those that you showed me ?

OpenStudy (anonymous):

sorry I made a mistake it should be 0.5² when n=3, 0.5^3 when n=4.... etc and when n=12 it should be 0.5^11

OpenStudy (anonymous):

Maybe it's a good idea for you to solve it like that once, so you know what you need to do when they ask you to calculate such a summation.

OpenStudy (tkhunny):

Yes. You should know how to add a finite geometric series. This WILL be required of you on exams.

OpenStudy (anonymous):

Give me a sec pretty please

OpenStudy (anonymous):

Can you help me please. I am clueless as to what ti do

OpenStudy (anonymous):

I Got 1409/2048 but i do not think that is correct

OpenStudy (anonymous):

or is it 20.54?

OpenStudy (anonymous):

\[20*0.5^{2}+20*0.5^{3}+20*0.5^{4}+20*0.5^{5}+20*0.5^{6}+20*0.5^{7}\] \[+20*0.5^{8}+20*0.5^{9}+20*0.5^{10}+20*0.5^{11}\] is what you need to calculate.

OpenStudy (anonymous):

Give me a sec

OpenStudy (anonymous):

How do you fan someone ?

OpenStudy (tkhunny):

Just look at it and say to yourself, "Blech!! I NEVER want to do that. There MUST be another way." Indeed, there is: \(5\cdot\left(1 + 1/2 + 1/4 + ... + 1/512 \right) = 5\cdot\dfrac{1 - \dfrac{1}{1024}}{1 - \dfrac{1}{2}}\) Thus, we have turned 9 additions, 10 multiplications, and 10 exponentiations into 3 divisions, 2 subtractions and one (1) multiplication. Just counting operations, we've come down from 29 to 6. Way, WAY less to go wrong. Way, WAY fewer rounding problems. Learn how to create that summation on the fly. This is the salient point of this exercise.